1.145 g of sodium chloride and 1.285 x 10^-2 mol of sodium peroxydisulfate were mixed together and ground in a mortar and pestle. Calculate the percent by mass of sodium chloride in the resulting mixture.
Answer #1

Molar mass of sodium peroxydisulfate (Na2S2O8) is 238.10 g/ mol
so we have (238.10 g / mol x 1.285x10-2 mol) = 3.060g sodium peroxydisulfate

total mass of the mixture = 3.060g + 1.145 g = 4.205g

% NaCl = 1.145 g / 4.205g = 27.2%